# Uncertainty Quantification on Graph Learning: A Survey

### v2

@@ VI Uncertainty Quantification Evaluation
VI-A Calibration
The evaluation of calibration measures the difference between a model’s prediction confidence and the exact probability that the predicted output is correct. This section introduces two widely applied metrics.
Expected Calibration Error (ECE). We divide the entire confidence space [0,1] into $M$ equal-size intervals $I_{m}=(\frac{m-1}{M},\frac{m}{M}]$ for $m=1,...,M$. $B_{m}$ is the set of samples with confidences in $I_{m}$. The accuracy and confidence of $B_{m}$ are
@@ VI-A Calibration
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum\nolimits_{i\in B_{m}}\hat{p}_{i}.$ (37)
respectively, where $y_{i}$ and $\hat{y}_{i}$ are the true and predicted label of input $x_{i}$, and $\hat{p}_{i}$ is the confidence of the prediction. Expected Calibration Error (ECE) computes the absolute difference between $\text{acc}(B_{m})$ and $\text{conf}(B_{m})$:
@@
To evaluate the calibration of node-level classification models on GNNs, node-wise ECE follows the concept of ECE. Specifically, test nodes are partitioned into bins according to their prediction confidences, and nodewise-ECE works as Eq. (38) to quantify the summation of the absolute difference between confidence and accuracy across all bins [135, 138, 134, 137]. However, nodewise-ECE neglects graph structure and the interdependencies between nodes. To remedy this, edge-wise ECE is proposed in [160]. With graph $G=(\mathcal{N},\mathcal{E})$, denote $\mathcal{E}_{U}=\{(i,j)\in\mathcal{E}|i,j\in U\}$ the subset of edges $\mathcal{E}$ constrained to test nodes $U$. The confidence of edge $(i,j)$ is $\hat{c}{(i,j)}=\max_{l,m}\hat{p}(y_{i}=l,y_{j}=m)$, given the features of all nodes and labels of training nodes. The calculation of edge-wise ECE is similar to node-wise ones. Specifically, each $\hat{c}{(i,j)}$ is computed and assigned to corresponding bins, and a prediction is considered accurate only if both predicted $l$ and $m$ are correct. The summed gap between confidence bins and accuracy is edge-wise ECE, similar to Eq. (38). Additionally, ECE for agreeing node pairs and disagreeing node pairs are also proposed [160].
Proper Scoring Rules. As ECE depends on the resolution of the finite number of bins, a continuous proper scoring rule, such as Brier Score (BS) [161], can act as an alternative to ECE. Brier Score (BS) considers a sample’s confidence in all labels. It is given by
$\text{BS}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{K}(\hat{p}_{i,j}-\mathbbm{1}(y_{i}=j)),$ (39)
where $\hat{p}_{i,j}$ is the confidence of sample $i$ in label $j$. A lower Brier Score indicates not only that a model is well-calibrated but also that the model is confident in its predictions.
VI-B Conformal Prediction
Conditional Coverage Metric. The key contribution of conformal prediction is stated in Eq. (29), requiring the coverage of prediction sets. However, it is not explicitly claimed but desired that a conformal prediction model can be adaptive for different conditions. For instance, we want the CP procedure can distinguish the difficulty of quantifying the prediction uncertainty of each input, and adaptively generate a larger prediction set as the difficulty increases [113]. In other words, the $1-\alpha$ coverage can be robust and stable no matter how the test domain changes. Let $E_{1},...,E_{N}$ denote the $N$ sets of test inputs coming from the corresponding test domains, and we want the coverage guarantee to always hold that $\forall e\in\{E_{1},...,E_{N}\}$:
@@ VI-B Conformal Prediction
This test domain can be characterized by an explicit feature in input data or an underlying environment variable related to the data generation process. Fig. 2 shows the differences between marginal and conditional coverage with $0.81-\alpha=$ 0.8. The worst coverage among multiple domains is given by
@@
where $\mathcal{I}_{e}$ represent the sample index set of set $e$ and $|\mathcal{I}_{e}|$ denote the number of samples in set $e$ [162, 163]. If the metric value is much lower than the expected coverage $1-\alpha$, we say the CP procedure performs poorly on conditional coverage.
Coverage Correctness. The empirical coverage by split conformal prediction involves randomness due to the limited number of available samples and possibly biased data splitting. Running conformal prediction multiple times can help us determine the correctness and stability of prediction coverage. With $n$ labeled samples available, we can randomly split them into $n_{cal}$ samples for calibration and $n_{val}$ samples for validation, and calculate the coverage by
@@
Figure 2: Marginal Coverage vs Conditional Coverage with (1-$\alpha$)=0.8. Marginal coverage can only guarantee the coverage on the overall sample set, but conditional coverage can ensure at least 1-$\alpha$ coverage on different sub-domians
VII Resources
This section compiles tutorials from prominent machine-learning conferences and open-source codes from some cited papers on related topics for convenient access.
TABLE III: Open-source Codes
@@ VII Resources
TABLE IV: Tutorials on Uncertainty Quantification and Graphic Models
@@ A Primer on PAC-Bayesian Learning ICML 2019
VIII Conclusion and Future Work
In this survey, we carefully reviewed the uncertainty quantification within graphical models, specifically focusing on Graph Neural Networks and Probabilistic Graphical Models. We examed existing methods for representing and handling uncertainty, including Bayesian representation learning, conformal prediction, calibration and so on. The domain of uncertainty in graphical models presents substantial opportunities for further innovation and advancement. Advancing our understanding of uncertainty modeling in graph-based systems will empower machine learning systems to make more reliable decisions in the presence of real-world noise. Such progress will not only improve the accuracy and safety of AI applications but also promote trust and encourage the broader adoption of these technologies in practical settings.

### v3

@@ VI Uncertainty Quantification Evaluation
This section discusses the evaluation of uncertainty quantification, including methods for direct inference and OOD data, calibration and conformal prediction.
VI-A Evaluating uncertainty due to OOD data
Unlike calibration and conformal prediction approaches that explicitly quantify and optimize predictive uncertainty, DRO, AT and SSL methods primarily target improved generalization under distributional shifts, e.g., OOD scenarios or adversarial attacks, without directly modeling uncertainty. In particular, DRO typically assesses model robustness by evaluating performance across shifted data distributions, while AT focuses on enhancing resilience to adversarial perturbations, evaluating both clean and adversarial predictive performance. Consequently, the evaluation of these techniques through the lens of uncertainty remains underexplored, highlighting a critical need for rigorous uncertainty-centric assessment frameworks.
VI-B Calibration
The evaluation of calibration measures the difference between a model’s prediction confidence and the exact probability that the predicted output is correct. Two widely applied metrics are introduced below.
Expected Calibration Error (ECE). The entire confidence space [0,1] can be divided into $M$ equal-size intervals $I_{m}=(\frac{m-1}{M},\frac{m}{M}]$ for $m=1,...,M$. $B_{m}$ is the set of samples with confidences in $I_{m}$. The accuracy and confidence of $B_{m}$ are
@@ VI-B Calibration
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum\nolimits_{i\in B_{m}}\hat{p}_{i},$ (43)
where $y_{i}$ and $\hat{y}_{i}$ are the true and predicted label of input $x_{i}$, and $\hat{p}_{i}$ is the confidence of the prediction. Expected Calibration Error (ECE) computes the absolute difference between $\text{acc}(B_{m})$ and $\text{conf}(B_{m})$:
@@
To evaluate the calibration of node-level classification models on GNNs, node-wise ECE follows the concept of ECE. Specifically, test nodes are partitioned into bins according to their prediction confidences, and nodewise-ECE works as Eq. (44) to quantify the summation of the absolute difference between confidence and accuracy across all bins [99, 102, 195, 101]. However, nodewise-ECE neglects graph structure and the interdependencies between nodes. To remedy this, edge-wise ECE is proposed in [226]. With graph $G=(\mathcal{V},\mathcal{E})$, denote $\mathcal{E}_{U}=\{(i,j)\in\mathcal{E}|i,j\in U\}$ the subset of edges $\mathcal{E}$ constrained to test nodes $U$. Given a trained model, the confidence of edge $(i,j)\in\mathcal{E}_{U}$ is $\hat{c}{(i,j)}=\hat{p}_{i}\hat{p}_{j}$. The calculation of edge-wise ECE is similar to node-wise ones. Specifically, each $\hat{c}{(i,j)}$ is computed and assigned to corresponding bins, and a prediction is considered accurate only if both predictions are correct (i.e. $y_{i}=\hat{y}_{i},y_{j}=\hat{y}_{j}$). The summed gap between confidence bins and accuracy is edge-wise ECE, similar to Eq. (44). Additionally, ECE for agreeing node pairs and disagreeing node pairs are also proposed [226].
Brier Score. As ECE depends on the resolution of the finite number of bins, a continuous proper scoring rule, such as Brier Score (BS) [227], can act as an alternative to ECE. BS considers a sample’s confidence in all labels. It is given by
$\text{BS}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{K}(v_{i}^{j}-\mathbbm{1}(y_{i}=j)),$ (45)
where $v_{i}^{j}$ is the probability that sample $i$ is labeled $j$ and $\mathbbm{1}(\cdot)$ is indicator function. A lower Brier Score indicates not only that a model is well-calibrated but also that the model is confident in its predictions.
VI-C Conformal Prediction
Conditional Coverage Metric. The key contribution of CP is the guaranteed coverage of prediction sets, as stated in Eq. (40). However, it is not explicitly claimed but desired that CP can be adaptive for different inputs (i.e. larger prediction sets for more uncertain predictions) [209]. In other words, the $1-\alpha$ coverage can be robust and stable no matter how the test domain changes. Let $E_{1},...,E_{N}$ denote the $N$ sets of test inputs coming from the corresponding test domains, and we want the coverage guarantee to always hold that $\forall e\in\{E_{1},...,E_{N}\}$:
@@ VI-C Conformal Prediction
This test domain can be characterized by an explicit feature in input data or an underlying environment variable related to the data generation process. The worst coverage among multiple domains is given by
@@
where $\mathcal{I}_{e}$ represent the sample index set of set $e$ and $|\mathcal{I}_{e}|$ denote the number of samples in set $e$ [228, 229]. If the metric value is much lower than the expected coverage $1-\alpha$, the CP procedure performs poorly on conditional coverage.
Coverage Correctness. The empirical coverage by split conformal prediction involves randomness due to the limited number of available samples and possibly biased data splitting. Running conformal prediction multiple times can help determine the prediction coverage’s correctness and stability. With $n$ labeled samples available, we can randomly split them into $n_{cal}$ samples for calibration and $n_{val}$ samples for validation, and calculate the coverage by
@@
VII Future Directions
In this section, we analyze existing challenges in uncertainty quantification on graphs, and summarize some future research directions facing the challenges.
Bayesian learning can generate credible regions for model predictions, but these regions are only valid if the model is correctly specified, meaning that the prior, likelihood, and computational capability must be accurately defined during posterior inference [230, 231, 232]. [230] primarily addresses these issues by introducing a temperature parameter into Bayesian GNNs within the conformal prediction framework. The combination of Bayesian inference with the conformal prediction framework allows for both credible regions and prediction sets that are more robust, even in the presence of model misspecification. Integrating Bayesian methods with other uncertainty quantification techniques could further mitigate inaccuracies arising from violations of Bayesian learning assumptions. By leveraging the strengths of complementary uncertainty quantification methods, such integrations could lead to more accurate and reliable uncertainty estimates for graph learning.
Out-of-distribution tasks have been well explored by various techniques, such as DRO, AT, and SSL. As discussed before, the majority of existing research has primarily focused on improving model generalization to OOD samples. However, there has been limited exploration into the deeper connections between UQ and OOD. A promising direction is to leverage OOD methods to better estimate and quantify uncertainties. Integrating OOD techniques with UQ could yield valuable insights, potentially leading to more robust and reliable uncertainty estimations. The integration could also help bridge the gap between OOD and UQ.
Conformal prediction for dynamic graphs presents a unique challenge. Most existing conformal prediction methods for graph data concentrate on static graphs within transductive and inductive frameworks to adhere to the exchangeability assumption or fixed distribution shift. However, dynamic graphs, such as those found in social networks, are common in real-world applications. These can result in dynamic distribution shifts, posing greater challenges for CP [233]. Addressing this challenge could enhance the interpretability of GNNs when dealing with evolving data, like changes in edges over time.
Calibration methods primarily address miscalibration errors by leveraging the homophily property of graphs, which assumes that neighboring nodes tend to belong to the same class [100, 99]. This approach has proven effective in many scenarios, but it falls short in heterophilous environments where neighboring nodes are more likely to belong to different classes. Despite the growing importance of heterophilous graphs, the calibration of GNNs in these contexts remains under-explored. Effectively calibrating GNNs under heterophily is essential for improving their reliability and performance in real-world applications characterized by complex relationships.
VIII Conclusion
In this survey, we carefully reviewed the uncertainty quantification within graphical models, specifically focusing on Graph Neural Networks and Probabilistic Graphical Models. We examined existing methods for representing and handling uncertainty, including Bayesian representation learning, Gaussian Processes, Outlier Representation, approaches based on Stochastic Differential Equations (SDEs), conformal prediction, calibration, and so on. The domain of uncertainty in graphical models presents substantial opportunities for further innovation and advancement. Advancing our understanding of uncertainty modeling in graph-based systems will empower machine learning systems to make more reliable decisions in the presence of real-world noise. Such progress will not only improve the accuracy and safety of AI applications but also promote trust and encourage the broader adoption of these technologies in practical settings.
Appendix A Resources
For convenient access, Table IV and Table V summarize open-source code repositories from some cited papers on related topics and tutorials from prominent machine learning conferences, respectively.
TABLE IV: Repositories of Open-source Codes
@@ Appendix A Resources
TABLE V: Tutorials on Uncertainty Quantification and Graphic Models